### v3

@@ Abstract.
The weak-cop number of a graph, a variation of the cop number, is an invariant suitable for infinite graphs and is a quasi-isometric invariant. While for any $m\in\mathbb{Z}_{+}\cup\{\infty\}$ there exist locally finite infinite graphs with weak-cop number $m$, it is an open question whether there exists locally finite vertex transitive graphs whose weak-cop number is different than $1$ and $\infty$. We test this question on Cayley graphs of wreath products, these are objects known for their exotic geometries. We prove that Cayley graphs of wreath products of nontrivial groups by infinite groups have infinite weak-cop number. The result is proved by defining a new pursuit and evasion game and proving the existence of strategies for the evader. We also include a short argument that Cayley graphs of Thompson’s group $F$ have infinite weak cop number.
@@ 1. Introduction
Lee et al. [LMPRQ23] introduced a variation of the cops and robber game called Weak-Cops and Robbers. This is a two-player game where one of the players controls a finite set of cops, while the other controls a single robber. The objective of the cops is to protect arbitrarily large finite subgraphs of the underlying graph, subject to some parameters. These parameters are chosen by the players. First, the cops player chooses the number of cops, as well as the cops’ speed and reach. The robber player, knowing this information, chooses his speed and challenges the cops to protect a large ball in the graph; the cops choose their initial positions, and then the robber choose his initial position. After these choices have been made, the game starts and the cops and robber move in alternating turns, up to a distance determined by their respective speeds. The cops win the game if at some stage the robber is captured (the robber is within reach of a cop) or from some turn on, they keep the robber outside the ball. See Section  2 for a precise definition.
The weak-cop number of a connected graph $\Gamma$, denoted $\operatorname{wCop}(\Gamma)$, is the minimum number of cops required for the cop player to always have a winning strategy in the Weak-Cops and Robbers game on $\Gamma$. If no such finite number exists, then we say that $\operatorname{wCop}(\Gamma)=\infty$. This is an interesting numerical invariant of infinite graphs, it is preserved by quasi-isometries, for every $m\in\mathbb{N}\cup\{\infty\}$ there exists a graph with weak-cop number $m$, and the invariant behaves well with respect to different operations including products, see the work of Lee et al [LMPRQ23]. They raised the following question given that their examples of graphs with arbitrary weak cop numbers are locally finite but lack symmetry.
@@
A natural source to explore this question is on Cayley graphs of finitely generated groups. In fact, the notion of weak-cop number yields an invariant of finitely generated groups, via playing the game on Cayley Graphs with respect to finite generating sets, since any two Cayley graphs with respect to finite generating sets of a group have the same weak-cop number [LMPRQ23, Corollary G].
Definition 1.2 (Weak-cop number of a group).
Let $G$ be a finitely generated group. The weak-cop number $\operatorname{wCop}(G)$ is defined as the weak-cop number of any Cayley graph of $G$ with respect to a finite generating set.
In [LMPRQ23], Lee et al showed that for finitely generated groups, free groups have weak-cop number 1, and non-cyclic free abelian groups have infinite weak cop number. A non-trivial result in their article is that one-ended non-amenable groups have infinite weak-cop number [LMPRQ23, Theorem H]. They also stated they following restricted version of Question 1.1.
@@
Question 1.3 is intriguing. A positive answer would place the weak-cop number as a rich integral invariant of finitely generated groups worth of a deeper study. Even in the negative case, a description of the class of finitely generated groups with weak-cop number one does not seem trivial, but it is plausible to expect a positive answer to the following question.
@@
For finitely generated groups, does having weak-cop number one implies being virtually free?
@@
The restricted wreath product $G\wr H$ of finitely generated groups has infinite weak-cop number if $G$ is non-trivial and $H$ is infinite.
@@
Theorem 1.5 is proved by defining a new game called Lamplighter. Briefly, this is a perfect information game with at least $P\geq 2$ players on a single underlying object called a streetmap, essentially a Cayley graph $\operatorname{Cay}(H)$ where the vertices represent lamps that can take on a variety of states given by $\operatorname{Cay}(G)$. One player, the lamplighter, moves along a copy of the graph $\operatorname{Cay}(H)$ restricted to a chosen finite subgraph called the area of play, changing the states of lamps. The other $P-1$ players, the copiers, move along their own copies of $\operatorname{Cay}(H)$, working together in an attempt for one one of them to approximate the lamplighter’s pattern of lamps. The lamplighters and the group of copiers move alternately, changing the state of some of the lamps close to their current position. The copiers win the game if eventually one of them can approximate the lamplighter’s pattern of lamps. Under the assumptions, we show that the lamplighter has a wining strategy for any number of copiers. We then prove that this implies that $\operatorname{Cay}(G\wr H)$ has infinite weak-cop number. This new game can be regarded as a variation of the weak-cops and robbers where the robber player imposes on himself a finite area of play, instead of being able to freely move on the underlying infinite graph. A precise definition of the Lamplighter game and the proof of Theorem 1.5 are the contents of Section 3.
Let us state a question related to Theorem 1.5. The Diestel-Leader graphs $DL(m,n)$ for $m,n\in\mathbb{Z}_{+}$ are connected, locally finite, vertex-transitive graphs which are regarded as generalizations of Cayley graphs of Lamplighter groups $\mathbb{Z}_{n}\wr\mathbb{Z}$. It is a remarkable result of Eskin, Fisher and Whyte that $D(m,n)$ is quasi-isometric to a Cayley graph of a finitely generated group if and only if $m=n$; see [EFW12]. The graph $D(m,m)$ is quasi-isometric to a Cayley graph of the Lamplighter group $\mathbb{Z}_{m}\wr\mathbb{Z}$ which has infinite weak-cop number by Theorem 1.5.
@@
Does $D(m,n)$ have infinite weak-cop number for any $m,n\in\mathbb{Z}_{+}$
Another finitely generated group known for its exotic geometry is Thompson’s group $F$, for background see [BG84, CFP96]. It is an outstanding question in group theory whether $F$ is amenable. In this note, we also include short computation of the weak-cop number of Thompson’s group $F$ communicated to the authors by Francesco Fournier-Facio.
@@
Thompson’s group $F$ has infinite weak-cop number.
@@
The rest of the article is organized into three sections. The first section contains some preliminaries and the precise definition of the weak-cop number for graphs. Then Section 3 contains the proof of Theorem 1.5. The last section contains the proof that Thompson’s group $F$ has infinite weak cop number.
@@ Acknowledgements
The authors thank Francesco Fournier-Facio, Florian Lehner and Danny Dyer for comments on preliminary versions of this work. The first author acknowledges funding by the Natural Sciences and Engineering Research Council of Canada NSERC, via the Undergraduate Student Research Award (USRA). The second author acknowledges funding by the Natural Sciences and Engineering Research Council of Canada NSERC.
@@ 2.1. Graph theory language
A graph $\Gamma$ is a pair $(V,E)$, where $V$ is called the set of vertices of $\Gamma$, and $E\subset\binom{V}{2}$ contains subsets of cardinality 2 in $V$, called the edges of $\Gamma$. We denote by $V(\Gamma)$ and $E(\Gamma)$ the vertex set and the edge set of $\Gamma$, respectively. Two vertices $u,v\in V$ are said to be adjacent in $\Gamma$ if $\{u,v\}\in E$. Observe that this definition encompasses simple graphs, i.e. graphs that have no edges from a vertex to itself (no loops), and each edge appears at most once in $E$ (no multiple edges). A graph is trivial if it has only one vertex, and is infinite if it has infinite vertex set. By an isomorphism from a graph $G$ to a graph $H$, we mean a bijection $\Phi\colon V(G)\rightarrow V(H)$ such that $\{u,v\}\in E(G)$ if and only if $\left\{\Phi(u),\Phi(v)\right\}\in E(H)$.
By a path in a graph $\Gamma$ we mean a sequence of vertices $v_{0}$, $v_{1}$, $v_{2}$, …, $v_{k}$ such that, for each $i=1,2,\ldots,k$, $\{v_{i},v_{i-1}\}$ is an edge in $\Gamma$. The length of such path is defined as $k$. A graph is connected if there if there is a path between any two vertices. In a connected graph, the length of the shortest path between two vertices $u,v$ is called the distance between them and is denoted by $\operatorname{\mathsf{dist}}_{\Gamma}(u,v)$. A path $v_{0},v_{1},\ldots,v_{k}$ is a geodesic if $\operatorname{\mathsf{dist}}_{\Gamma}(v_{0},v_{k})=k$. Note that in a connected graph, there is a geodesic between any pair of points. The diameter of $\Gamma$ is the supremum of the distances between any pair of vertices, in particular, a graph can have infinite diameter. Note that if the graph has infinite diameter, then there are geodesics of arbitrarily large length. A graph $\Gamma$ is locally finite if for any vertex $v$, the set of vertices at distance one from $v$ is finite.
@@
2.2. Definition of the weak-cop number
Given a connected graph $\Gamma$, Weak-Cops and Robbers is played on $\Gamma$ as follows. There are two players, with one playing the robber, and one playing a set of $n$ cops for some $n\in\mathbb{N}$. Before the game begins, the cops choose two positive integers $\sigma$ and $\rho$, called the cop’s speed and reach, respectively. Knowing these values, the robbers then choose positive integers $\psi$ and $R$, along with a vertex $v$ of $\Gamma$; these parameters are called the robber’s speed, the radius of the area of play and the center of the area of play respectively.
Once the parameters have been chosen, each of the cops choose a vertex for their initial positions. Knowing the choices made by each of the cops, the robber chooses a vertex for their own initial position. The cops and robber move in alternating turns, starting with the cops. On the cops’ turn, each one can move to a vertex at distance at most $\sigma$ from their current position. The robber is captured during this turn if any of the cops move to within distance $\rho$ of him. On the robber’s turn, he can move to a vertex at distance at most $\psi$ from his current position, provided he has a path to that vertex that contains no vertex within distance $\rho$ from any of the cops. The cops win if they can eventually protect the ball of radius $R$ centered at $v$. This means that either the robber is captured or, beginning on some turn, they can permanently prevent the robber from moving within distance $R$ of the vertex $v$.
@@ 2.2. Definition of the weak-cop number
The weak-cop number of $\Gamma$, denoted $\operatorname{wCop}(\Gamma)$, is the smallest $n$ such that $\Gamma$ is $n$-weak cop win. If no such integer exists, we say $\operatorname{wCop}(\Gamma)=\infty$.
3. The weak-cop number of wreath products
@@ 3.1.2. Turns and player movement
After all of the parameters are selected, play happens in turns, starting with the copiers. On the copiers’ turn, they all play in unison. The copiers are able to communicate with one another and coordinate their moves. On each turn, every player has a number of possible moves, and they choose a number of moves to play up to their speed. Let $B=(M,p,\phi)$ be the player’s board on a given turn. There are two types of moves, each one changing the player’s board:
@@
The copiers, on each of their turns, move up to a total of $\sigma$ times (being able to select any of their available moves each time), followed by the lamplighter similarly moving up to $\psi$ times.
@@ 3.1.4. Win conditions
Given a streetmap $M$, we denote by $\mathcal{B}(M)$ the set of all possible boards on $M$. The distance between two boards in $\mathcal{B}(M)$ is defined as the minimum number of moves required to change one board into the other.
The goal of the copiers is to come ”close enough” to copying the lamplighter’s board. Specifically, the copiers win if at any point during either player’s turn, the lamplighter’s board is at distance $\rho$ or less from any of the copiers’ boards. Conversely, the lamplighter wins if he can devise a strategy that will let him avoid coming within the copiers’ reach indefinitely.
@@ 3.2. General winning strategy for the Lamplighter
Note that, for each copier $c_{i}$, the argument above only relies on the position of $c_{i}$ and the states of lights $\ell_{i}$ and $r_{i}$. In particular, the position of the lamplighter does not affect this fact, nor do the states of the other lamps. Therefore, changing the lamplighter’s position does not cause the copiers to win. Additionally, this means that changing the state of a lamp $\ell_{i}$ or $r_{i}$ does not bring any copier $c_{j}$ with $j\neq i$ within distance $\rho$ of the lamplighter; in fact, it strictly increases the distance from the lamplighter to copier $c_{i}$. Thus, the copiers do not win during the lamplighter’s turn.
@@ 3.3. The wreath product of graphs
3.4. Relation Between Lamplighter and Weak-cops and Robbers
In order to use our Lamplighter game to derive results about Weak-cops and robbers, we must establish some additional theory.
@@ 3.4. Relation Between Lamplighter and Weak-cops and Robbers
There is a straightforward connection between the Lamplighter game and the Weak-cops and robbers game, described as follows.
@@
Let $\Lambda$ be a connected graph with infinite diameter and $\Omega$ a nontrivial connected graph. If $M=(\Omega,\omega,\Lambda)$ is a streetmap and $n$ a positive integer, then the winning strategy for the lamplighter given by Theorem 3.6 for the game of Lamplighter on $M$ with $n$ copiers, provides a winning strategy for the robber in a game of Weak-Cops and Robbers on $\Omega\wr\Lambda$ with $n$ cops.
@@
It follows that moving in a game of Lamplighter on $M$ (either changing positions or changing states) is equivalent to moving between vertices in Weak-cops and robbers on $\Omega\wr\Lambda$. Consequently, the lamplighter speed, copier reach, and copier speed in the Lamplighter game correspond exactly with the robber speed, reach, and cop speed, respectively in the Weak-cops and robbers game. The only parameters that do not align exactly are the radius of the area of play and the center of the area of play. According to the winning strategy for the lamplighter given by Theorem 3.6 via Remark 3.7, we have that $r=\left\lceil\frac{\sigma+\rho}{2}\right\rceil+n$ is the radius of the area of play, and the center of the area of play $v\in V(\Lambda)$ is at the center of a geodesic path $P$ in $\Lambda$ with $12r+$ 1 vertices. Moreover, if $(M,p_{1},\phi_{1})$ and $(M,p_{2},\phi_{2})$ are boards representing the lamplighter during distinct stages in the game then the distance between them is bounded from above by $R=2(6r+1)$.
The center of the area of play in the Weak-Cops and Robbers game is defined to be the board $(M,v,0)$ where 00 denotes the constant function $V(\Lambda)\to\{w\}$ and the radius of the area of play is defined as $R$. Note that Remark 3.7 states that any board representing the lamplighter is at distance at most $R$ from $(M,v,0)$.
Since the vertices of $\Omega\wr\Lambda$ have been identified with boards, the strategy for the robber in the Weak-cops and robbers game is defined as our winning strategy for the lamplighter against $n$ copiers that reproduce the moves of the cops in the Weak-Cops and Robbers game. Then, since at any stage of the Lamplighter game, the board representing the lamplighter is at distance larger than $\rho$ than any board representing a copier, we have that the robber is never captured. We show in the previous paragraph that any stage of the game, the board representing the lamplighter (and hence the robber) is in the $R$-ball about the center of play $(M,v,0)$ in $\Omega\wr\Lambda$ the Weak-cops and robbers game. Thus, this constitutes a winning strategy for the robber. ∎
The above results gives us the following corollary that shows that the weak-cop number of a larger class of wreath products of graphs is infinite.
@@
3.5. Weak-cop number of wreath-products of groups

### v4

@@ Abstract.
The weak cop number of a graph, a variation of the cop number, is an invariant suitable for infinite graphs and is a quasi-isometric invariant. While for any $m\in\mathbb{Z}_{+}\cup\{\infty\}$ there exist locally finite infinite graphs with weak cop number $m$, it is an open question whether there exists locally finite vertex transitive graphs whose weak cop number is different than $1$ and $\infty$. We test this question on Cayley graphs of wreath products; these are objects known for their exotic geometries. We prove that Cayley graphs of wreath products of nontrivial groups by infinite groups have infinite weak cop number. The result is proved by defining a new pursuit and evasion game and proving the existence of strategies for the evader. We also include a short argument that Cayley graphs of Thompson’s group $F$ have infinite weak cop number.
@@ 1. Introduction
Lee et al. [LMPRQ23] introduced a variation of the cops and robber game called Weak cops and Robbers. This is a two-player game where one of the players controls a finite set of cops, while the other controls a single robber. The objective of the cops is to protect arbitrarily large finite subgraphs of the underlying graph, subject to some parameters. These parameters are chosen by the players. First, the cops player chooses the number of cops, as well as the cops’ speed and reach. The robber player, knowing this information, chooses his speed and challenges the cops to protect a large ball in the graph; the cops choose their initial positions, and then the robber choose his initial position. After these choices have been made, the game starts and the cops and robber move in alternating turns, up to a distance determined by their respective speeds. The cops win the game if at some stage the robber is captured (the robber is within reach of a cop) or from some turn on, they keep the robber outside the ball. See Section  2 for a precise definition.
The weak cop number of a connected graph $\Gamma$, denoted $\operatorname{wCop}(\Gamma)$, is the minimum number of cops required for the cop player to always have a winning strategy in the Weak cops and Robbers game on $\Gamma$. If no such finite number exists, then we say that $\operatorname{wCop}(\Gamma)=\infty$. This is an interesting numerical invariant of infinite graphs, it is preserved by quasi-isometries, for every $m\in\mathbb{N}\cup\{\infty\}$ there exists a graph with weak cop number $m$, and the invariant behaves well with respect to different operations including products, see the work of Lee et al [LMPRQ23]. They raised the following question given that their examples of graphs with arbitrary weak cop numbers are locally finite but lack symmetry.
@@
A natural source to explore this question is on Cayley graphs of finitely generated groups. In fact, the notion of weak cop number yields an invariant of finitely generated groups, via playing the game on Cayley Graphs with respect to finite generating sets, since any two Cayley graphs with respect to finite generating sets of a common group have the same weak cop number [LMPRQ23, Corollary G].
Definition 1.2 (Weak cop number of a group).
Let $G$ be a finitely generated group. The weak cop number $\operatorname{wCop}(G)$ is defined as the weak cop number of any Cayley graph of $G$ with respect to a finite generating set.
In [LMPRQ23], Lee et al showed that for finitely generated groups, free groups have weak cop number 1, and non-cyclic free abelian groups have infinite weak cop number. A non-trivial result in their article is that one-ended non-amenable groups have infinite weak cop number [LMPRQ23, Theorem H]. They also stated they following restricted version of Question 1.1.
@@
Question 1.3 is intriguing. A positive answer would place the weak cop number as a rich integral invariant of finitely generated groups worth of a deeper study. Even in the negative case, a description of the class of finitely generated groups with weak cop number one does not seem trivial, but it is plausible to expect a positive answer to the following question.
@@
For finitely generated groups, does having weak cop number one implies being virtually free?
@@
The restricted wreath product $G\wr H$ of finitely generated groups has infinite weak cop number if $G$ is non-trivial and $H$ is infinite.
@@
Theorem 1.5 is proved by defining a new game called Lamplighter. Briefly, this is a perfect information game with at least $P\geq 2$ players on a single underlying object called a streetmap, essentially a Cayley graph $\operatorname{Cay}(H)$ where the vertices represent lamps that can take on a variety of states given by $\operatorname{Cay}(G)$. One player, the lamplighter, moves along a copy of the graph $\operatorname{Cay}(H)$ restricted to a chosen finite subgraph called the area of play, changing the states of lamps. The other $P-1$ players, the copiers, move along their own copies of $\operatorname{Cay}(H)$, working together in an attempt for one of them to approximate the lamplighter’s pattern of lamps. The lamplighter and the group of copiers move alternately, changing the state of some of the lamps close to their current position. The copiers win the game if eventually one of them can approximate the lamplighter’s pattern of lamps. Under the assumptions, we show that the lamplighter has a wining strategy for any number of copiers. We then prove that this implies that $\operatorname{Cay}(G\wr H)$ has infinite weak cop number. This new game can be regarded as a variation of the weak cops and robbers where the robber player imposes on himself a finite area of play, instead of being able to freely move on the underlying infinite graph. A precise definition of the Lamplighter game and the proof of Theorem 1.5 are the contents of Section 3.
Let us state a question related to Theorem 1.5. The Diestel-Leader graphs $DL(m,n)$ for $m,n\in\mathbb{Z}_{+}$ are connected, locally finite, vertex-transitive graphs which are regarded as generalizations of Cayley graphs of Lamplighter groups $\mathbb{Z}_{n}\wr\mathbb{Z}$. It is a remarkable result of Eskin, Fisher and Whyte that $D(m,n)$ is quasi-isometric to a Cayley graph of a finitely generated group if and only if $m=n$; see [EFW12]. The graph $D(m,m)$ is quasi-isometric to a Cayley graph of the Lamplighter group $\mathbb{Z}_{m}\wr\mathbb{Z}$ which has infinite weak cop number by Theorem 1.5.
@@
Does $D(m,n)$ have infinite weak cop number for any $m,n\in\mathbb{Z}_{+}$
Another finitely generated group known for its exotic geometry is Thompson’s group $F$, for background see [BG84, CFP96]. It is an outstanding question in group theory whether $F$ is amenable. In this note, we also include short computation of the weak cop number of Thompson’s group $F$ communicated to the authors by Francesco Fournier-Facio.
@@
Thompson’s group $F$ has infinite weak cop number.
@@
The rest of the article is organized into three sections. Section 2 contains some preliminaries and the precise definition of the weak cop number for graphs. Then Section 3 contains the proof of Theorem 1.5. The last section contains the proof that Thompson’s group $F$ has infinite weak cop number.
@@ Acknowledgements
The authors thank Francesco Fournier-Facio, Florian Lehner and Danny Dyer for comments on preliminary versions of this work. Special thanks to the anonymous referee for suggestions and corrections. The first author acknowledges funding by the Natural Sciences and Engineering Research Council of Canada NSERC, via the Undergraduate Student Research Award (USRA). The second author acknowledges funding by the Natural Sciences and Engineering Research Council of Canada NSERC.
@@ 2.1. Graph theory language
A graph $\Gamma$ is a pair $(V,E)$, where $V$ is called the set of vertices of $\Gamma$, and $E\subset\binom{V}{2}$ contains subsets of cardinality 2 in $V$, called the edges of $\Gamma$. Let $V(\Gamma)$ and $E(\Gamma)$ denote the vertex set and the edge set of $\Gamma$ respectively. Two vertices $u,v\in V$ are said to be adjacent in $\Gamma$ if $\{u,v\}\in E$. Observe that this definition encompasses simple graphs, i.e. graphs that have no edges from a vertex to itself (no loops), and each edge appears at most once in $E$ (no multiple edges). A graph is trivial if it has only one vertex, and is infinite if it has infinite vertex set. By an isomorphism from a graph $G$ to a graph $H$, we mean a bijection $\Phi\colon V(G)\rightarrow V(H)$ such that $\{u,v\}\in E(G)$ if and only if $\left\{\Phi(u),\Phi(v)\right\}\in E(H)$.
By a path in a graph $\Gamma$ we mean a sequence of vertices $v_{0}$, $v_{1}$, $v_{2}$, …, $v_{k}$ such that, for each $i=1,2,\ldots,k$, $\{v_{i},v_{i-1}\}$ is an edge in $\Gamma$. The length of such path is defined as $k$. A graph is connected if there if there is a path between any two vertices. In a connected graph, the length of the shortest path between two vertices $u,v$ is called the distance between them and is denoted by $\operatorname{\mathsf{dist}}_{\Gamma}(u,v)$. A path $v_{0},v_{1},\ldots,v_{k}$ is a geodesic if $\operatorname{\mathsf{dist}}_{\Gamma}(v_{0},v_{k})=k$. Note that in a connected graph, there is a geodesic between any pair of points. The diameter of $\Gamma$ is the supremum of the distances between any pair of vertices; in particular, a graph can have infinite diameter. Note that if the graph has infinite diameter, then there are geodesics of arbitrarily large length. A graph $\Gamma$ is locally finite if for any vertex $v$, the set of vertices at distance one from $v$ is finite.
@@
2.2. Definition of the weak cop number
Given a connected graph $\Gamma$, Weak Cops and Robbers is played on $\Gamma$ as follows. There are two players, with one playing the robber, and one playing a set of $n$ cops for some $n\in\mathbb{N}$. Before the game begins, the cops choose two positive integers $\sigma$ and $\rho$, called the cops’ speed and reach, respectively. Knowing these values, the robber then chooses positive integers $\psi$ and $R$, along with a vertex $v$ of $\Gamma$; these parameters are called the robber’s speed, the radius of the area of play and the center of the area of play respectively.
Once the parameters have been chosen, each of the cops choose a vertex for their initial positions. Knowing the choices made by each of the cops, the robber chooses a vertex for their own initial position. The cops and robber move in alternating turns, starting with the cops. On the cops’ turn, each cop can move to a vertex at distance at most $\sigma$ from its current position. The robber is captured during this turn if any of the cops move to within distance $\rho$ of him. On the robber’s turn, he can move to a vertex at distance at most $\psi$ from his current position, provided he has a path to that vertex that contains no vertex within distance $\rho$ from any of the cops. The cops win if they can eventually protect the ball of radius $R$ centered at $v$. This means that either the robber is captured or, beginning on some turn, they can permanently prevent the robber from moving within distance $R$ of the vertex $v$.
@@ 2.2. Definition of the weak cop number
The weak cop number of $\Gamma$, denoted $\operatorname{wCop}(\Gamma)$, is the smallest $n$ such that $\Gamma$ is $n$-weak cop win. If no such integer exists, we say $\operatorname{wCop}(\Gamma)=\infty$.
3. The weak cop number of wreath products
@@ 3.1.2. Turns and player movement
After all of the parameters are selected, play happens in turns, starting with the copiers. On the copiers’ turn, they all play in unison. The copiers are able to communicate with one another and coordinate their moves. The copiers, on each of their turns, move up to a total of $\sigma$ times (being able to select any of their available moves each time), followed by the lamplighter similarly moving up to $\psi$ times.
Let $B=(M,p,\phi)$ be the player’s board on a given turn. There are two types of moves, each one changing the player’s board:
@@ 3.1.4. Win conditions
Given a streetmap $M$, we denote by $\mathcal{B}(M)$ the set of all possible boards on $M$. Define the distance between two boards in $\mathcal{B}(M)$ to be the minimum number of moves required to change one board into the other. Note that this defines a metric on $\mathcal{B}(M)$.
The goal of the copiers is to come “close enough” to copying the lamplighter’s board. Specifically, the copiers win if at any point during either player’s turn, the lamplighter’s board is at distance $\rho$ or less from any of the copiers’ boards. Conversely, the lamplighter wins if he can devise a strategy that will let him avoid coming within the copiers’ reach indefinitely.
@@ 3.2. General winning strategy for the Lamplighter
Note that, for each copier $c_{i}$, the argument above only relies on the position of $c_{i}$ and the states of lights $\ell_{i}$ and $r_{i}$. In particular, the position of the lamplighter does not affect this fact, nor do the states of the other lamps. Therefore, changing the lamplighter’s position does not cause the copiers to win. Additionally, this means that changing the state of a lamp $\ell_{i}$ or $r_{i}$ does not bring any copier $c_{j}$ with $j\neq i$ within distance $\rho$ of the lamplighter; furthermore, it strictly increases the distance from the lamplighter to copier $c_{i}$. Thus, the copiers do not win during the lamplighter’s turn.
@@ 3.3. The wreath product of graphs
3.4. Relation Between Lamplighter and Weak cops and Robbers
In order to use our Lamplighter game to derive results about Weak cops and robbers, we must establish some additional theory.
@@ 3.4. Relation Between Lamplighter and Weak cops and Robbers
There is a straightforward connection between the Lamplighter game and the Weak cops and robbers game, described as follows.
@@
Let $\Lambda$ be a connected graph with infinite diameter and $\Omega$ a nontrivial connected graph. If $M=(\Omega,\omega,\Lambda)$ is a streetmap and $n$ a positive integer, then the winning strategy for the lamplighter given by Theorem 3.6 for the game of Lamplighter on $M$ with $n$ copiers, provides a winning strategy for the robber in a game of Weak Cops and Robbers on $\Omega\wr\Lambda$ with $n$ cops.
@@
It follows that moving in a game of Lamplighter on $M$ (either changing positions or changing states) is equivalent to moving between vertices in Weak cops and robbers on $\Omega\wr\Lambda$. Consequently, the lamplighter speed, copier reach, and copier speed in the Lamplighter game correspond exactly with the robber speed, reach, and cop speed, respectively in the Weak Cops and Robbers game. The only parameters that do not align exactly are the radius of the area of play and the center of the area of play. According to the winning strategy for the lamplighter given by Theorem 3.6 via Remark 3.7, we have that $r=\left\lceil\frac{\sigma+\rho}{2}\right\rceil+n$ is the radius of the area of play, and the center of the area of play $v\in V(\Lambda)$ is at the center of a geodesic path $P$ in $\Lambda$ with $12r+$ 1 vertices. Moreover, if $(M,p_{1},\phi_{1})$ and $(M,p_{2},\phi_{2})$ are boards representing the lamplighter during distinct stages in the game then the distance between them is bounded from above by $R=2(6r+1)$.
The center of the area of play in the Weak Cops and Robbers game is defined to be the board $(M,v,0)$ where 00 denotes the constant function $V(\Lambda)\to\{w\}$ and the radius of the area of play is defined as $R$. Note that Remark 3.7 states that any board representing the lamplighter is at distance at most $R$ from $(M,v,0)$.
Since the vertices of $\Omega\wr\Lambda$ have been identified with boards, the strategy for the robber in the Weak Cops and robbers game is defined as our winning strategy for the lamplighter against $n$ copiers that reproduce the moves of the cops in the Weak Cops and Robbers game. Then, since at any stage of the Lamplighter game, the board representing the lamplighter is at distance larger than $\rho$ than any board representing a copier, we have that the robber is never captured. We show in the previous paragraph that any stage of the game, the board representing the lamplighter (and hence the robber) is in the $R$-ball about the center of play $(M,v,0)$ in $\Omega\wr\Lambda$ the Weak Cops and Robbers game. Thus, this constitutes a winning strategy for the robber. ∎
The above results gives us the following corollary that shows that the weak cop number of a larger class of wreath products of graphs is infinite.
@@
3.5. Weak cop number of wreath-products of groups